# Jacobian Descent For Multi-Objective Optimization

### v1

@@ B.3 DualProj
The projection of a gradient of interest onto a dual cone was first described in [33]. When this gradient is the average of the rows of the Jacobian, we call this aggregator $\mathcal{A}_{\text{DualProj}}$. Formally, $\mathcal{A}_{\text{DualProj}}(J)=\frac{1}{m}\cdot\pi_{J}(J^{\top}\cdot{\boldsymbol{1}})$. By construction, $\mathcal{A}_{\text{DualProj}}$ is thus non-conflicting. By Proposition 1, $\mathcal{A}_{\text{DualProj}}(J)=\frac{1}{m}J^{\top}\cdot{\boldsymbol{w}}$, with ${\boldsymbol{w}}\in\operatorname*{arg\,min}_{{\boldsymbol{1}}\preceq{\boldsymbol{v}}}{\boldsymbol{v}}^{\top}JJ^{\top}{\boldsymbol{v}}$. $\mathcal{A}_{\text{DualProj}}$ is thus weighted.
@@ B.4 PCGrad
$\mathcal{A}_{\text{PCGrad}}$ is described in [11]. It projects each gradient onto the orthogonal hyperplane of other gradients in case of conflict with them, iteratively and in random order. When $m\varleq 2$, $\mathcal{A}_{\text{PCGrad}}$ is deterministic and is such that $\mathcal{A}_{\text{PCGrad}}=m\cdot\mathcal{A}_{\text{UPGrad}}$. Therefore, in this case, it satisfies all three properties. When $m>2$, $\mathcal{A}_{\text{PCGrad}}$ is non-deterministic, so $\mathcal{A}_{\text{PCGrad}}(J)$ is a random vector.
@@
For any index $i\in[m]$, let ${\boldsymbol{g}}_{i}=J^{\top}\cdot{\boldsymbol{e}}_{i}$ and let $P(i)$ be distributed uniformly on the permutations of the elements in $[m]\setminus\{i\}$. For instance, if $m=3$, $P(2)=\begin{bmatrix}1&3\end{bmatrix}^{\top}$ with probability $0.5$ and $P(2)=\begin{bmatrix}3&1\end{bmatrix}^{\top}$ with probability $0.5$. The iterative projection of $\mathcal{A}_{\text{PCGrad}}$ is then defined recursively as:
${\boldsymbol{g}}_{i,1}^{\text{PC}}$ $={\boldsymbol{g}}_{i}$
$j_{k}$ $=P(i)_{k}$
${\boldsymbol{g}}_{i,k+1}^{\text{PC}}$ $={\boldsymbol{g}}_{i,k}^{\text{PC}}-\mathds{1}\{{\boldsymbol{g}}_{i,k}^{\text{PC}}\cdot{\boldsymbol{g}}_{j_{k}}<0\}\frac{{\boldsymbol{g}}_{i,k}^{\text{PC}}\cdot{\boldsymbol{g}}_{j_{k}}}{\|{\boldsymbol{g}}_{j_{k}}\|^{2}}{\boldsymbol{g}}_{j_{k}}$ (20)
@@
We noticed that an equivalent formulation to the conditional projection of (20) is the projection onto the dual cone of $\{{\boldsymbol{g}}_{j_{k}}\}$:
${\boldsymbol{g}}_{i,k+1}^{\text{PC}}=\pi_{{\boldsymbol{g}}_{j_{k}}^{\top}}({\boldsymbol{g}}_{i,k}^{\text{PC}})$ (21)
Finally, the aggregation is given by $\mathcal{A}_{\text{PCGrad}}(J)=\sum_{i=1}^{m}{\boldsymbol{g}}_{i,m}^{\text{PC}}$.
For all $i$, ${\boldsymbol{g}}_{i,m}^{\text{PC}}$ is always a linear combination of rows of $J$. $\mathcal{A}_{\text{PCGrad}}$ is thus weighted.
@@
If $J=\begin{bmatrix}1&0\\ 0&1\\ -0.5&-1\end{bmatrix}$, the only non-conflicting direction is ${\boldsymbol{0}}$. However, $\mathcal{A}_{\text{PCGrad}}(J)$ is uniform over the set $\left\{\begin{bmatrix}0.4\\ 0.2\end{bmatrix},\begin{bmatrix}0.8\\ 0.2\end{bmatrix},\begin{bmatrix}0.4\\ -0.2\end{bmatrix},\begin{bmatrix}0.8\\ -0.2\end{bmatrix}\right\}$. $\mathcal{A}_{\text{PCGrad}}$ is thus not non-conflicting. Note that here, $\mathds{E}[\mathcal{A}_{\text{PCGrad}}(J)]=\begin{bmatrix}0.6&0\end{bmatrix}^{\top}$, so $\mathcal{A}_{\text{PCGrad}}$ is neither non-conflicting in expectation.
@@
To show that $\mathcal{A}_{\text{PCGrad}}$ is linear under scaling, let ${\boldsymbol{0}}\prec{\boldsymbol{c}}\in{\mathbb{R}}^{m}$, ${\boldsymbol{g}}_{i}^{\prime}=c_{i}{\boldsymbol{g}}_{i}$, ${\boldsymbol{g}}_{i,1}^{\prime\text{ PC}}={\boldsymbol{g}}^{\prime}_{i}$ and ${\boldsymbol{g}}_{i,k+1}^{\prime\text{ PC}}=\pi_{{\boldsymbol{g}}_{j_{k}}^{\prime\top}}({\boldsymbol{g}}_{i,k}^{\prime\text{ PC}})$. We show by induction that ${\boldsymbol{g}}_{i,k}^{\prime\text{ PC}}=c_{i}{\boldsymbol{g}}_{i,k}^{\text{PC}}$.
The base case is given by ${\boldsymbol{g}}_{i,1}^{\prime\text{ PC}}={\boldsymbol{g}}_{i}^{\prime}=c_{i}{\boldsymbol{g}}_{i}=c_{i}{\boldsymbol{g}}_{i,1}^{\text{PC}}$.
Then, assuming the induction hypothesis ${\boldsymbol{g}}_{i,k}^{\prime\text{ PC}}=c_{i}{\boldsymbol{g}}_{i,k}^{\text{PC}}$, we show ${\boldsymbol{g}}_{i,k+1}^{\prime\text{ PC}}=c_{i}{\boldsymbol{g}}_{i,k+1}^{\text{PC}}$:
@@
${\boldsymbol{g}}_{i,k+1}^{\prime\text{ PC}}=\pi_{c_{j_{k}}{\boldsymbol{g}}_{j_{k}}^{\top}}(c_{i}{\boldsymbol{g}}_{i,k}^{\text{ PC}})$ (Induction $\left(\begin{subarray}{c}\text{Induction hypothesis}\end{subarray}\right)$
${\boldsymbol{g}}_{i,k+1}^{\prime\text{ PC}}=c_{i}\pi_{{\boldsymbol{g}}_{j_{k}}^{\top}}({\boldsymbol{g}}_{i,k}^{\text{ PC}})$ $\left(\begin{subarray}{c}0<c_{i}\text{ and }0<c_{j_{k}}\end{subarray}\right)$
@@
Therefore $\mathcal{A}_{\text{PCGrad}}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)=\sum_{i=1}^{m}c_{i}{\boldsymbol{g}}_{i,m}^{\text{PC}}$, so $\mathcal{A}_{\text{PCGrad}}$ is linear under scaling.
@@ B.5 GradDrop
The aggregator used by the GradDrop layer, which we denote $\mathcal{A}_{\text{GradDrop}}$, is described in [17]. It is non-deterministic, so $\mathcal{A}_{\text{GradDrop}}(J)$ is a random vector. Given $J\in\mathbb{R}^{m\times n}$, let $|J|\in\mathbb{R}^{m\times n}$ be the element-wise absolute value of $J$. Let $P\in{\mathbb{R}}^{n}$ be such that $P=\frac{1}{2}\left({\boldsymbol{1}}+\frac{J^{\top}\cdot{\boldsymbol{1}}}{|J|^{\top}\cdot{\boldsymbol{1}}}\right)$, where the division is element-wise. Each coordinate $i\in[n]$ is independently assigned to the set $A$ with probability $P_{i}$ and to the set $B$ otherwise. The aggregation at coordinate $i\in A$ is given by the sum of all positive $J_{ji}$, for $j\in[m]$. The aggregation at coordinate $i\in B$ is given by the sum of all negative $J_{ji}$, for $j\in[m]$.
If $J=\begin{bmatrix}-2\\ 1\end{bmatrix}$, then $P=\begin{bmatrix}1/3\end{bmatrix}$. Therefore, $\mathds{P}[\mathcal{A}_{\text{GradDrop}}(J)=\begin{bmatrix}-2\end{bmatrix}]=2/3$ and $\mathds{P}[\mathcal{A}_{\text{GradDrop}}(J)=\begin{bmatrix}1\end{bmatrix}]=1/3$. Therefore, $\mathcal{A}_{\text{GradDrop}}$ is not non-conflicting (even in expectation).
$\mathcal{A}_{\text{GradDrop}}$ is not linear under scaling. Indeed, suppose that $J=\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$ then $P=\frac{1}{2}\cdot{\boldsymbol{1}}$ and the aggregation is one of the four vectors [±$\begin{bmatrix}\pm 1&\pm 1\end{bmatrix}^{\top}$ with equal probability. Scaling the first line of $J$ by $2$ yields $J=\begin{bmatrix}2&-2\\ -1&1\end{bmatrix}$, $P=\begin{bmatrix}2/3&1/3\end{bmatrix}^{\top}$ which cannot lead to a uniform distribution over four elements.
@@ B.6 IMTL-G
In [13], the authors describe a method to impartially balance gradients by weighting them. Let ${\boldsymbol{g}}_{i}$ be the $i$’th row of $J$ and let ${\boldsymbol{u}}_{i}=\frac{{\boldsymbol{g}}_{i}}{\|{\boldsymbol{g}}_{i}\|}$. Then, they want to find a combination ${\boldsymbol{g}}=\sum_{i=1}^{m}\alpha_{i}{\boldsymbol{g}}_{i}$ such that ${\boldsymbol{g}}^{\top}{\boldsymbol{u}}_{i}$ is equal for all $i$. Let $U=\begin{bmatrix}{\boldsymbol{u}}_{1}-{\boldsymbol{u}}_{2}&\dots&{\boldsymbol{u}}_{1}-{\boldsymbol{u}}_{m}\end{bmatrix}^{\top}$, $D=\begin{bmatrix}{\boldsymbol{g}}_{1}-{\boldsymbol{g}}_{2}&\dots&{\boldsymbol{g}}_{1}-{\boldsymbol{g}}_{m}\end{bmatrix}^{\top}$. If ${\boldsymbol{\alpha}}_{2:m}=\begin{bmatrix}\alpha_{2}&\dots&\alpha_{m}\end{bmatrix}^{\top}$, then ${\boldsymbol{\alpha}}_{2:m}=(UD^{\top})^{-1}U\cdot{\boldsymbol{g}}_{1}$ and $\alpha_{1}=1-\sum_{i=2}^{m}\alpha_{i}$. Notice that this is defined only when the gradients are linearly independent. Strictly speaking, this is thus not an aggregator since it can only be computed on matrices of rank $m$. To generalize to the case where $J$ has rank less than $m$, we provide what we believe to be a more natural formulation, which is equivalent in the case where $J$ has rank $m$. Let $J\in\mathbb{R}^{m\times n}$ and $J^{\prime}\in\mathbb{R}^{m\times n}$ be the row normalized version of $J$. Reformulating the desired property that ${\boldsymbol{g}}^{\top}{\boldsymbol{u}}_{i}$ is constant over $i$, in our framework, we solve for the following objective. Let ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ be such that ${\boldsymbol{1}}^{\top}{\boldsymbol{w}}=1$ and $J^{\prime}J^{\top}{\boldsymbol{w}}~{}\propto~{}{\boldsymbol{1}}$. The aggregated vector is then $J^{\top}{\boldsymbol{w}}$. If $J^{\prime}J^{\top}$ is full rank, then ${\boldsymbol{w}}~{}\propto~{}(J^{\prime}J^{\top})^{-1}{\boldsymbol{1}}$. Letting ${\boldsymbol{d}}\in{\mathbb{R}}^{m}$ be the vector of norms of the rows of $J$, this can be rewritten as ${\boldsymbol{w}}~{}\propto~{}(JJ^{\top})^{-1}{\boldsymbol{d}}$. When $JJ^{\top}$ is not full rank, using the Moore-Penrose inverse instead of the usual inverse seems to be a reasonable generalization. The vector of weights is then given by ${\boldsymbol{w}}~{}\propto~{}(JJ^{\top})^{\dagger}{\boldsymbol{d}}$. Finally whenever ${\boldsymbol{1}}^{\top}(JJ^{\top})^{\dagger}{\boldsymbol{d}}=0$, set ${\boldsymbol{w}}={\boldsymbol{0}}$. Throughout this paper, $\mathcal{A}_{\text{IMTL-G}}$ refers to this generalization.
@@
Clearly, $\mathcal{A}_{\text{IMTL-G}}$ is weighted. We now show that $\mathcal{A}_{\text{IMTL-G}}$ is not non-conflicting. If $J=\begin{bmatrix}1&-1&-1\end{bmatrix}^{\top}$, then ${\boldsymbol{d}}=\begin{bmatrix}1&1&1\end{bmatrix}^{\top}$. Also $JJ^{\top}=\begin{bmatrix}1&-1&-1\\ -1&1&1\\ -1&1&1\end{bmatrix}$ and therefore $(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}=\frac{1}{9}\begin{bmatrix}-1&1&1\end{bmatrix}^{\top}$. Note that in order to ensure ${\boldsymbol{1}}^{\top}{\boldsymbol{w}}=1$, we have to scale by ${\boldsymbol{1}}^{\top}(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}=\frac{1}{9}$. Therefore ${\boldsymbol{w}}=\begin{bmatrix}-1&1&1\end{bmatrix}^{\top}$, $\mathcal{A}_{\text{IMTL-G}}(J)=J^{\top}{\boldsymbol{w}}=\begin{bmatrix}-3\end{bmatrix}$ and $J\cdot\mathcal{A}_{\text{IMTL-G}}(J)=\begin{bmatrix}-3&3&3\end{bmatrix}^{\top}$. $\mathcal{A}_{\text{IMTL-G}}$ is thus not non-conflicting.
@@
It should be noted that when $J$ has rank $m$, $\mathcal{A}_{\text{IMTL-G}}$ seems to be non-conflicting. Thus, it would be possible to make a different non-conflicting generalization, for instance, by deciding ${\boldsymbol{0}}$ when the matrix is not full rank.
$\mathcal{A}_{\text{IMTL-G}}$ is not linear under scaling. if ${\boldsymbol{0}}\prec{\boldsymbol{c}}\in{\mathbb{R}}^{m}$, then the weights associated to $\operatorname*{diag}({\boldsymbol{c}})\cdot J$ are proportional to $(\operatorname*{diag}({\boldsymbol{c}})\cdot JJ^{\top}\cdot\operatorname*{diag}({\boldsymbol{c}}))^{\dagger}\cdot\operatorname*{diag}({\boldsymbol{c}})\cdot{\boldsymbol{d}}=\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}$. The normalization value is therefore $\frac{1}{{\boldsymbol{1}}^{\top}\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}$. Therefore
$\mathcal{A}_{\text{IMTL-G}}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)$ $=\frac{J^{\top}\cdot\operatorname*{diag}({\boldsymbol{c}})\cdot\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}{{\boldsymbol{1}}^{\top}\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}$
$=\frac{J^{\top}(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}{{\boldsymbol{1}}^{\top}\operatorname*{diag}({\boldsymbol{c}})^{\dagger}\cdot(JJ^{\top})^{\dagger}\cdot{\boldsymbol{d}}}$
@@
This is not linear in ${\boldsymbol{c}}$, so $\mathcal{A}_{\text{IMTL-G}}$ is not linear under scaling.

### v2

@@ B.3 DualProj
The projection of a gradient of interest onto a dual cone was first described in Lopez-Paz & Ranzato (2017). When this gradient is the average of the rows of the Jacobian, we call this aggregator $\mathcal{A}_{\text{DualProj}}$. Formally,
$\mathcal{A}_{\text{DualProj}}(J)=\frac{1}{m}\cdot\pi_{J}\left(J^{\top}\cdot{\bm{1}}\right)$ (22)
where $\pi_{J}$ is the projection operator defined in (3).
  Non-conflicting.
By the constraint in (3), $\mathcal{A}_{\text{DualProj}}$ is non-conflicting.
  Linear under scaling.
@@
  Weighted.
By Proposition 1, $\mathcal{A}_{\text{DualProj}}(J)=\frac{1}{m}J^{\top}\cdot{\bm{w}}$, with ${\bm{w}}\in\operatorname*{arg\,min}_{{\bm{1}}\preceq{\bm{v}}}{\bm{v}}^{\top}JJ^{\top}{\bm{v}}$. $\mathcal{A}_{\text{DualProj}}$ is thus weighted.
@@ B.4 PCGrad
$\mathcal{A}_{\text{PCGrad}}$ is described in Yu et al. (2020). It projects each gradient onto the orthogonal hyperplane of other gradients in case of conflict with them, iteratively and in random order. When $m\varleq 2$, $\mathcal{A}_{\text{PCGrad}}$ is deterministic and satisfies $\mathcal{A}_{\text{PCGrad}}=m\cdot\mathcal{A}_{\text{UPGrad}}$. Therefore, in this case, it satisfies all three properties. When $m>2$, $\mathcal{A}_{\text{PCGrad}}$ is non-deterministic, so $\mathcal{A}_{\text{PCGrad}}(J)$ is a random vector.
@@
For any index $i\in[m]$, let ${\bm{g}}_{i}=J^{\top}\cdot{\bm{e}_{i}}$ and let ${\mathbf{p}}(i)$ be a random vector distributed uniformly on the set of permutations of the elements in $[m]\setminus\{i\}$. For instance, if $m=3$, ${\mathbf{p}}(2)=\begin{bmatrix}1&3\end{bmatrix}^{\top}$ with probability $0.5$ and ${\mathbf{p}}(2)=\begin{bmatrix}3&1\end{bmatrix}^{\top}$ with probability $0.5$. For notation convenience, whenever $i$ is clear from context, we denote $j_{k}={\mathbf{p}}(i)_{k}$. The iterative projection of $\mathcal{A}_{\text{PCGrad}}$ is then defined recursively as:
${\bm{g}}_{i,1}^{\text{PC}}$ $={\bm{g}}_{i}$ (23)
${\bm{g}}_{i,k+1}^{\text{PC}}$ $={\bm{g}}_{i,k}^{\text{PC}}-\mathds{1}\{{\bm{g}}_{i,k}^{\text{PC}}\cdot{\bm{g}}_{j_{k}}<0\}\frac{{\bm{g}}_{i,k}^{\text{PC}}\cdot{\bm{g}}_{j_{k}}}{\|{\bm{g}}_{j_{k}}\|^{2}}{\bm{g}}_{j_{k}}$ (24)
@@
We noticed that an equivalent formulation to the conditional projection of (24) is the projection onto the dual cone of $\{{\bm{g}}_{j_{k}}\}$:
${\bm{g}}_{i,k+1}^{\text{PC}}=\pi_{{\bm{g}}_{j_{k}}^{\top}}({\bm{g}}_{i,k}^{\text{PC}})$ (25)
Finally, the aggregation is given by
$\mathcal{A}_{\text{PCGrad}}(J)=\sum_{i=1}^{m}{\bm{g}}_{i,m}^{\text{PC}}.$ (26)
  Non-conflicting.
@@
If $J=\begin{bmatrix}1&0\\ 0&1\\ -0.5&-1\end{bmatrix}$, the only non-conflicting direction is ${\bm{0}}$. However, $\mathcal{A}_{\text{PCGrad}}(J)$ is uniform over the set $\left\{\begin{bmatrix}0.4\\ 0.2\end{bmatrix},\begin{bmatrix}0.8\\ 0.2\end{bmatrix},\begin{bmatrix}0.4\\ -0.2\end{bmatrix},\begin{bmatrix}0.8\\ -0.2\end{bmatrix}\right\}$, i.e. $\mathcal{A}_{\text{PCGrad}}(J)$ is in the dual cone of the rows of $J$ with probability 00. $\mathcal{A}_{\text{PCGrad}}$ is thus not non-conflicting. Here, $\mathds{E}[\mathcal{A}_{\text{PCGrad}}(J)]=\begin{bmatrix}0.6&0\end{bmatrix}^{\top}$, so $\mathcal{A}_{\text{PCGrad}}$ is neither non-conflicting in expectation.
  Linear under scaling.
@@
To show that $\mathcal{A}_{\text{PCGrad}}$ is linear under scaling, let ${\bm{0}}\prec{\bm{c}}\in{\mathbb{R}}^{m}$, ${\bm{g}}_{i}^{\prime}=c_{i}{\bm{g}}_{i}$, ${\bm{g}}_{i,1}^{\prime\text{ PC}}={\bm{g}}^{\prime}_{i}$ and ${\bm{g}}_{i,k+1}^{\prime\text{ PC}}=\pi_{{\bm{g}}_{j_{k}}^{\prime\top}}\big{(}{\bm{g}}_{i,k}^{\prime\text{ PC}}\big{)}$. We show by induction that ${\bm{g}}_{i,k}^{\prime\text{ PC}}=c_{i}{\bm{g}}_{i,k}^{\text{PC}}$.
The base case is given by ${\bm{g}}_{i,1}^{\prime\text{ PC}}={\bm{g}}_{i}^{\prime}=c_{i}{\bm{g}}_{i}=c_{i}{\bm{g}}_{i,1}^{\text{PC}}$.
Then, assuming the induction hypothesis ${\bm{g}}_{i,k}^{\prime\text{ PC}}=c_{i}{\bm{g}}_{i,k}^{\text{PC}}$, we show ${\bm{g}}_{i,k+1}^{\prime\text{ PC}}=c_{i}{\bm{g}}_{i,k+1}^{\text{PC}}$:
@@
${\bm{g}}_{i,k+1}^{\prime\text{ PC}}=\pi_{c_{j_{k}}{\bm{g}}_{j_{k}}^{\top}}\big{(}c_{i}{\bm{g}}_{i,k}^{\text{ PC}}\big{)}$ (Induction $\left(\begin{subarray}{c}\text{Induction hypothesis}\end{subarray}\right)$
${\bm{g}}_{i,k+1}^{\prime\text{ PC}}=c_{i}\pi_{{\bm{g}}_{j_{k}}^{\top}}\big{(}{\bm{g}}_{i,k}^{\text{ PC}}\big{)}$ $\left(\begin{subarray}{c}0<c_{i}\text{ and }0<c_{j_{k}}\end{subarray}\right)$
@@
Therefore $\mathcal{A}_{\text{PCGrad}}\big{(}\operatorname*{diag}({\bm{c}})\cdot J\big{)}=\sum_{i=1}^{m}c_{i}{\bm{g}}_{i,m}^{\text{PC}}$, so it can be written as $\mathcal{A}_{\text{PCGrad}}\big{(}\operatorname*{diag}({\bm{c}})\cdot J\big{)}=\mathfrak{J}^{\top}\cdot{\bm{c}}$ with $\mathfrak{J}=\begin{bmatrix}{\bm{g}}_{1,m}^{\text{PC}}&\cdots&{\bm{g}}_{m,m}^{\text{PC}}\end{bmatrix}^{\top}$. Therefore, $\mathcal{A}_{\text{PCGrad}}$ is linear under scaling.
  Weighted.
For all $i$, ${\bm{g}}_{i,m}^{\text{PC}}$ is always a random linear combination of rows of $J$. $\mathcal{A}_{\text{PCGrad}}$ is thus weighted.
@@ B.5 GradDrop
The aggregator used by the GradDrop layer, which we denote $\mathcal{A}_{\text{GradDrop}}$, is described in Chen et al. (2020). It is non-deterministic, so $\mathcal{A}_{\text{GradDrop}}(J)$ is a random vector. Given $J\in\mathbb{R}^{m\times n}$, let $|J|\in\mathbb{R}^{m\times n}$ be the element-wise absolute value of $J$. Let $P=\frac{1}{2}\left({\bm{1}}+\frac{J^{\top}\cdot{\bm{1}}}{|J|^{\top}\cdot{\bm{1}}}\right)\in{\mathbb{R}}^{n}$, where the division is element-wise. Each coordinate $i\in[n]$ is independently assigned to the set $\mathcal{I}_{+}$ with probability $P_{i}$ and to the set $\mathcal{I}_{-}$ otherwise. The aggregation at coordinate $i\in\mathcal{I}_{+}$ is given by the sum of all positive $J_{ji}$, for $j\in[m]$. The aggregation at coordinate $i\in\mathcal{I}_{-}$ is given by the sum of all negative $J_{ji}$, for $j\in[m]$. Formally,
$\mathcal{A}_{\text{GradDrop}}(J)$ $=\left(\sum_{i\in\mathcal{I}_{+}}{\bm{e}_{i}}\sum_{\begin{subarray}{c}j\in[m]:\\ J_{ji}>0\end{subarray}}J_{ji}\right)+\left(\sum_{i\in\mathcal{I}_{-}}{\bm{e}_{i}}\sum_{\begin{subarray}{c}j\in[m]:\\ J_{ji}<0\end{subarray}}J_{ji}\right)$ (27)
  Non-conflicting.
If $J=\begin{bmatrix}-2\\ 1\end{bmatrix}$, then $P=\begin{bmatrix}1/3\end{bmatrix}$. Therefore, $\mathds{P}\big{[}\mathcal{A}_{\text{GradDrop}}(J)=\begin{bmatrix}-2\end{bmatrix}\big{]}=2/3$ and $\mathds{P}\big{[}\mathcal{A}_{\text{GradDrop}}(J)=\begin{bmatrix}1\end{bmatrix}\big{]}=1/3$, i.e. $\mathcal{A}_{\text{GradDrop}}(J)$ is in the dual cone of the rows of $J$ with probability 00. Therefore, $\mathcal{A}_{\text{GradDrop}}$ is not non-conflicting. Here, $\mathds{E}\big{[}\mathcal{A}_{\text{GradDrop}}(J)\big{]}=\begin{bmatrix}-1\end{bmatrix}^{\top}$, so $\mathcal{A}_{\text{GradDrop}}$ is neither non-conflicting in expectation.
  Linear under scaling.
If $J=\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$, then $P=\frac{1}{2}\cdot{\bm{1}}$ and the aggregation is one of the four vectors [±$\begin{bmatrix}\pm 1&\pm 1\end{bmatrix}^{\top}$ with equal probability. Scaling the first line of $J$ by $2$ yields $J=\begin{bmatrix}2&-2\\ -1&1\end{bmatrix}$ and $P=\begin{bmatrix}2/3&1/3\end{bmatrix}^{\top}$, which cannot lead to a uniform distribution over four elements. Therefore, $\mathcal{A}_{\text{GradDrop}}$ is not linear under scaling.
  Weighted.
@@ B.6 IMTL-G
In Liu et al. (2021b), the authors describe a method to impartially balance gradients by weighting them. Let ${\bm{g}}_{i}$ be the $i$’th row of $J$ and let ${\bm{u}}_{i}=\frac{{\bm{g}}_{i}}{\|{\bm{g}}_{i}\|}$. They want to find a combination ${\bm{g}}=\sum_{i=1}^{m}\alpha_{i}{\bm{g}}_{i}$ such that ${\bm{g}}^{\top}{\bm{u}}_{i}$ is equal for all $i$. Let $U=\begin{bmatrix}{\bm{u}}_{1}-{\bm{u}}_{2}&\dots&{\bm{u}}_{1}-{\bm{u}}_{m}\end{bmatrix}^{\top}$, $D=\begin{bmatrix}{\bm{g}}_{1}-{\bm{g}}_{2}&\dots&{\bm{g}}_{1}-{\bm{g}}_{m}\end{bmatrix}^{\top}$. If ${\bm{\alpha}}_{2:m}=\begin{bmatrix}\alpha_{2}&\dots&\alpha_{m}\end{bmatrix}^{\top}$, then ${\bm{\alpha}}_{2:m}=\big{(}UD^{\top}\big{)}^{-1}U\cdot{\bm{g}}_{1}$ and $\alpha_{1}=1-\sum_{i=2}^{m}\alpha_{i}$. Notice that this is defined only when the gradients are linearly independent. Thus, this is not strictly speaking an aggregator since it can only be computed on matrices of rank $m$. We thus propose a generalization defined for matrices of any rank that is equivalent when the matrix has rank $m$. In the original formulation, requiring ${\bm{g}}^{\top}{\bm{u}}_{i}$ to be equal to some $c\in{\mathbb{R}}$ for all $i$, is equivalent to requiring that for all $i$, ${\bm{g}}^{\top}{\bm{g}}_{i}$ is equal to $c~{}\|{\bm{g}}_{i}\|$. Writing ${\bm{g}}=J^{\top}{\bm{\alpha}}$ and letting ${\bm{d}}\in{\mathbb{R}}^{m}$ be the vector of norms of the rows of $J$, the objective is thus to find ${\bm{\alpha}}$ satisfying $JJ^{\top}{\bm{\alpha}}\propto{\bm{d}}$. Besides, to match the original formulation, the elements of ${\bm{\alpha}}$ should sum to $1$.
Letting $\big{(}JJ^{\top}\big{)}^{\dagger}$ be the Moore-Penrose pseudo inverse of $JJ^{\top}$, we define
$\mathcal{A}_{\text{IMTL-G}}(J)$ $=J^{\top}\cdot{\bm{w}}$ (28)
$\text{with}\hskip 19.91684pt{\bm{w}}$ =$=\begin{cases}\frac{{\bm{v}}}{{\bm{1}}^{\top}{\bm{v}}},&\text{if }{\bm{1}}^{\top}{\bm{v}}\neq{\bm{0}}\\ {\bm{0}},&\text{otherwise}\end{cases}$ (29)
$\text{and}\hskip 25.03856pt{\bm{v}}$ $=\big{(}JJ^{\top}\big{)}^{\dagger}\cdot{\bm{d}}.$ (30)
  Non-conflicting.
@@
If $J=\begin{bmatrix}1&-1&-1\end{bmatrix}^{\top}$, then ${\bm{d}}=\begin{bmatrix}1&1&1\end{bmatrix}^{\top}$, $JJ^{\top}=\begin{bmatrix}1&-1&-1\\ -1&1&1\\ -1&1&1\end{bmatrix}$, and thus ${\bm{v}}=\frac{1}{9}\begin{bmatrix}-1&1&1\end{bmatrix}^{\top}$. Therefore, ${\bm{w}}=\begin{bmatrix}-1&1&1\end{bmatrix}^{\top}$, $\mathcal{A}_{\text{IMTL-G}}(J)=\begin{bmatrix}-3\end{bmatrix}^{\top}$ and $J\cdot\mathcal{A}_{\text{IMTL-G}}(J)=\begin{bmatrix}-3&3&3\end{bmatrix}^{\top}$. $\mathcal{A}_{\text{IMTL-G}}$ is thus not non-conflicting.
@@
It should be noted that when $J$ has rank $m$, $\mathcal{A}_{\text{IMTL-G}}$ seems to be non-conflicting. Thus, it would be possible to make a different non-conflicting generalization, for instance, by deciding ${\bm{0}}$ when $J$ is not full rank.
  Linear under scaling.
@@
With $J=\begin{bmatrix}a&0\\ 0&1\end{bmatrix}$ and $a>0$, we have ${\bm{v}}=\begin{bmatrix}1/a^{2}&0\\ 0&1\end{bmatrix}\cdot\begin{bmatrix}a\\ 1\end{bmatrix}=\begin{bmatrix}1/a\\ 1\end{bmatrix}$ and $\mathcal{A}_{\text{IMTL-G}}(J)=\begin{bmatrix}a&0\\ 0&1\end{bmatrix}\cdot\begin{bmatrix}1/a\\ 1\end{bmatrix}\cdot\frac{1}{\frac{1}{a}+1}=\frac{1}{\frac{1}{a}+1}\cdot\begin{bmatrix}1\\ 1\end{bmatrix}$. This is not affine in $a$, so $\mathcal{A}_{\text{IMTL-G}}$ is not linear under scaling.
  Weighted.
By (28), $\mathcal{A}_{\text{IMTL-G}}$ is weighted.